the perimeter of a rectangle is 12x+4. if the width is 2x, the length of the rectangle is....
Answer by stanbon(75887)

the perimeter of a rectangle is 12x+4. if the width is 2x, the length of the rectangle is....
--------------------------------------
Perimeter = 2(L + W)
12x+4 = 2(L + 2x)
6x + 2 = L + 2x
L = 4x + 2
===================
Cheers,
Stan H.